OpenStudy (anonymous):

How do you solve for x in this equation? 3 = log8 + 3logx

OpenStudy (anonymous):

By law of logs nloga => logaⁿ, we have: 3 = log(8) + log(x³) When adding logs, you are multiplying the terms altogether. 3 = log(8x³) Then, by log_a b = n ==> aⁿ = b: 10³ = 8x³ [Since the base is not given, I assume that the base is 10] 1000 = 8x³ Finally, solve for x. 1000/8 = x³ x³ = 125 x = (125)^(1/3) . . .Set both sides to the power of 1/3. x = 5 Hence, x = 5.

OpenStudy (anonymous):

x=5 does no appear to be the correct answer. Refer to the attached from WolframAlpha.

OpenStudy (anonymous):

Thank youuuu

OpenStudy (anonymous):

Hmm... Interesting @robtobey . Thanks for the heads up!

OpenStudy (anonymous):

Sorry, left a t off of the end of no.

OpenStudy (anonymous):

Me or tobey? @cayssaday cause Tobey got the right one

OpenStudy (anonymous):

Both of you !!:)

OpenStudy (anonymous):

Lol no prob! ;)

OpenStudy (anonymous):

Both of you put the thought into helping me so thank you :)